OpenStudy (joshoyen):

Write each function in vertex form and identify its vertex. 1. f(x) = x^2 + 6x -7 2. f(x) = 2x^2 - 12x - 27

OpenStudy (lochana):

vertex form is\[y= a(x-h)^2 + k\]

OpenStudy (lochana):

can you do it now?

OpenStudy (joshoyen):

uhm not sure

OpenStudy (joshoyen):

still stuck:/

OpenStudy (lochana):

okay. I will do first one\[y = x^2+6x -7 \\ y = (x^2 +6x + 9) -9 -7 (9\ is \ square\ of\ half \ of\ the\ coefficient\ of \ 'x')\]

OpenStudy (joshoyen):

hmm, wait how did you get 9 again? its the half of the coefficient of x, idk how that works

OpenStudy (lochana):

hey, I am back\[6x \rightarrow 6 \rightarrow \frac{6}{2} \rightarrow 3 \rightarrow 3^2 \rightarrow 9 \]

OpenStudy (lochana):

that's how 9 comes in

OpenStudy (lochana):

\[y = (x+3)^2 - 16 \\ vertex point (-3,-16)\]

OpenStudy (joshoyen):

Ohhhhhh, I see now, ok thank you so much